OpenStudy (anonymous):

The sum of Pete's and Sam's ages is 30. Five years ago, Pete was 3 times as old as Sam. How old is Sam? Let P = Pete's age, S = Sam's age, and P + S = 30. Which of the following equations would complete the system? P - 5 = 3S P - 5 = 3S - 5 P - 5 = 3S - 15

OpenStudy (tkhunny):

Please show your work. The sum of Pete's and Sam's ages is 30 P + S = 30 Five years ago, Pete was 3 times as old as Sam Translate this sentence.

OpenStudy (kylexchaos):

P + S = 30 since five years ago, Pete was /three times/ as old as Sam... P-5=3S-5 you have to subtract 5 from both sides, or else the equation would not be equal.